A rectangular field is 56m wide. It is three times as long as it is wide.What is its perimeter in metres?
Answer by Timnewman(323)

hi!
If it's width(w)=56m,
it's lenght =3w
=3*56
=168m
The perimeter (p)=2(L+w)
p=2(168+56)
p=2(224m)
p=448m
The perimetre of the rectangular field is 448m
BRAVO
---Timnewman